OpenStudy (anonymous):

find the second derivative in terms of x and y y^2 = x^3

OpenStudy (dumbcow):

find 1st derivative \[2y \frac{dy}{dx} = 3x^{2}\] \[\frac{dy}{dx} = \frac{3x^{2}}{2y}\] differentiate again using quotient rule \[\frac{d^{2}y}{dx} = \frac{(6x)(2y)-(3x^{2})(2dy/dx)}{(2y)^{2}}\] sub in 1st derivative and simplify